mikewwe13:

1. Refer to the check register and bank statement to answer the questions.

mikewwe13:

1 attachment
mikewwe13:

1 attachment
mikewwe13:

A. Paycheck 645.30 should be added to balance of 411.07  654.30 + 411.07 = 1065.37 B. Just add the two deposits, 654 + 654 = 1308 C. Add the total debits 1582.5 + 168.36 + 157.42 + 148.46 = 2056.74 D. Beginning balance - debits + deposits 3986.26 - 2056.74 + 1290 = 3219.52 E. Look for the negative balance....cable bill caused this -85 F. Check it to see, i'm not sure G. Statement period says it only goes through 8/3

mikewwe13:

am i right @jhonyy9

jhonyy9:

beginning balance = 398,26 now will be more easy than you assume first the so add to beginning balance the deposits and after this subtract from this what you get so the total sum of withdrawal

jhonyy9:

the last one calcule not is correct

mikewwe13:

398,26 + 1290 = 1688.26

mikewwe13:

Which of the letters are correct and not correct ?

jhonyy9:

the 8/06 line

jhonyy9:

you have not completed in this table the 7/31 line

jhonyy9:

too

mikewwe13:

645 + 30 = 675

jhonyy9:

no these 30 are cents

mikewwe13:

OH OKAY, IS IT 645 + 645 = 1290

jhonyy9:

in these second columns are cents

mikewwe13:

okay, i understand

jhonyy9:

but you dont need make the calcule newly just correct what i ve wrote above that not is correct

jhonyy9:

so step by step in the line of 7/31 you missed a sum from there

mikewwe13:

645 + 645 = 1290 1290 - 30 = 1260

jhonyy9:

no not this

jhonyy9:

check the line of 7/24 and the line of 7/31 in the line of 7/31 you dont assumed nothing in the BALANCE column

jhonyy9:

on this table where the last line is 8/06

jhonyy9:

bc. in one of this table this line of 8/06 dont exist you missed it from there

jhonyy9:

do you can more quickly please bc. here is in the night 0,12 oclock

mikewwe13:

i've been trying and i still don't know how to do it cause it's always wrong

mikewwe13:

sorry man , i'm trying to speed quickly as i can to get it done.

jhonyy9:

you forget assuming the 411,07 and 645,30 = ?

mikewwe13:

oh yes and i didn't know from the rest 411,07 + 645,30 = 1056.37

jhonyy9:

ok and now from this you need subtract this 80,00 from the date of 8/06 the last down line of this table so here was your mistake

jhonyy9:

???

mikewwe13:

1056.37 - 80,00 =

jhonyy9:

1056,37 -80 = ?

mikewwe13:

976.37

jhonyy9:

ok and look please what you wrote in this place of this table - what result ?

jhonyy9:

986,37 so what is correct ?

mikewwe13:

1056.37

jhonyy9:

no what is correct to ending balance

mikewwe13:

80

jhonyy9:

no i think you confused all here on the end

mikewwe13:

i think i am

mikewwe13:

was it 411,07

jhonyy9:

and ?

mikewwe13:

645,30

jhonyy9:

yes the last deposit

jhonyy9:

and ?

mikewwe13:

148.46

jhonyy9:

no you need assuming these two sum of

jhonyy9:

bc. 411,07 is the balance on what will be deposited the last deposit sum of 645,30

mikewwe13:

oh okay

mikewwe13:

it might be wrong , was it 1056.37

jhonyy9:

why is wrong ? not is sure so this sum will give to you 1056,37 from what you need subtract the last one withdrawal of 80,0

jhonyy9:

= ?

jhonyy9:

1056,37 - 80 = ?

mikewwe13:

976.37

mikewwe13:

i wanted to be sure cause i was afraid to be wrong

jhonyy9:

ok but can you tel me then why you wrote there in this table like last result 986,37 ?

jhonyy9:

so do you can complete now this table lines of total withdrawals total deposits and ending balance ?

mikewwe13:

because i was unsure and confused then and i think so

mikewwe13:

was A, B, C D, E , and G right ?

jhonyy9:

no just A,B,C,D,E and F not was completed and G was incorrect

mikewwe13:

so my answers on those was right and G was wrong ?

jhonyy9:

so do you can complete now this table lines of total withdrawals total deposits and ending balance ?

jhonyy9:

F not was completed and G was wrong

mikewwe13:

i don't think so i can complete it if i was shown to do it and i see

jhonyy9:

do you like i wait and check your work ?

mikewwe13:

yes

jhonyy9:

ok step by step how you get the total sum of deposits ?

mikewwe13:

by subtracting

jhonyy9:

so you need assuming these sums from deposit column

jhonyy9:

do you see this table with deposit column ?

jhonyy9:

there are two the same sums of 645,30 yes ? how much is the total of these ?

mikewwe13:

1290

jhonyy9:

and the cents

mikewwe13:

30

jhonyy9:

two times this 30

jhonyy9:

645,30 + 645,30 ???????

mikewwe13:

1290.6

jhonyy9:

,60

jhonyy9:

so this is the total deposited sum

jhonyy9:

and now the total withdrawal sum

jhonyy9:

check this table with column of withdrawals

mikewwe13:

80

jhonyy9:

and you need assuming these 5 sums of

jhonyy9:

80 is the last one

mikewwe13:

158, 168, 157, 148, 80

jhonyy9:

yes

mikewwe13:

do i add or subtract ?

jhonyy9:

assume what mean to you ?

mikewwe13:

yes, what comes next ?

jhonyy9:

add this sums

mikewwe13:

158 + 168 + 157 + 148 + 80 = 711

jhonyy9:

yes

jhonyy9:

no sorry where are these cents from there ?

jhonyy9:

in the second column are cents what you need adding too

mikewwe13:

Ohh sorry

mikewwe13:

it was on the left side

mikewwe13:

25 + 36 + 42 + 47 + 80 = 711

jhonyy9:

do you play with me now ?

jhonyy9:

25 ,36,42 is ok and what is there again on this column ?

jhonyy9:

46 and 00 so these are the cents what you need adding again too

mikewwe13:

ooh i see your game now

jhonyy9:

game ???

mikewwe13:

what you saying , i meant like in another expression when you said do you play with me now ... you get what i mean

mikewwe13:

46 + 00 = ?

jhonyy9:

and just now ???

jhonyy9:

yes the first 3 and these 46 +00 so total = ?

mikewwe13:

46

jhonyy9:

and these 3 first wrote before

jhonyy9:

so you need adding again these sum of cents from second column of these withdrawals sum and so will get the total withdrawals

jhonyy9:

the ending balance we ve calcule it firstly on the top these are all hope helped bye bye

mikewwe13:

so this is for G and F ?

mikewwe13:

could you point out which ones goes for those for me to put ?

mikewwe13:

thank you

mikewwe13:

thank you